A spherical balloon filled with helium gas must lift a1.00-kg scientific payload off the ground, and the mass of the empty balloon is 0.500 kg. The mass densities are 1.286 kg/m3 for air, 0.179 kg/m3for helium gas, 1.251 kg/m3 for nitrogen gas.

Part A

What is the minimum diameter of the balloon?

Express your answer with the appropriate units.

d =

SubmitMy AnswersGive Up

Part B

A helium shortage forces the scientists to consider using nitrogen gas in the balloon. What minimum balloon diameter is needed if nitrogen gas is used?

Express your answer with the appropriate units.

d =
Answer #1

Baur on - 0:5 Py em aur (1) a =(M+42)I e l-286V = 1.5や179 V -107 3 R= 0.686 m tam ittf 1342 m Draueter 434 m